355 in a sample of 1000 individuals said they would vote for candidate A in the next election. If candidate A obtained 33.2% (.332) of the vote in the last election, would you conclude that candidate A will obtain a higher percentage of the vote in the next election? Obtain the p-value Question options: .939 .061 .071 .081
Answer #1

H0: p = 0.332

Ha: p > 0.332

Sample proportion = 355 / 1000 = 0.355

Test statistics

z = ( - p) / sqrt [ p ( 1 - p) / n ]

= ( 0.355 - 0.332) / sqrt [ 0.332 ( 1 - 0.332) / 1000 ]

= 1.544

p-value = P(Z > z)

= P(Z > 1.544)

= 0.061 (From Z table)